Week of 190708

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Alberto (Monitoring), Aritz (Storage), Borja (Chair, Monitoring), Julien (Storage), Maarten (ALICE), Olga (Computing), Remi (Storage), Sabine (ATLAS)
  • remote: Andrew (NL-T1), Christoph (CMS), Darren (RAL), Di (TRIUMF), Marcelo (CNAF), Samuel (KIT), Sang-Un (KISTI), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Smooth production for jobs (~ 400k) and data (15 GB/s)
      • Heavy Ions overlay and small reprocessing campaigns started, overlay jobs caped to avoid too much stress on frontier servers
      • Benefiting from T0 and CERN-P1 however not running at 100%
    • Finalising new monitoring tests, old DDM monitoring switched off, old job monitoring to be switched off at the end of the week
    • Starting moving production queues to pilot2 + containers.

  • CMS reports ( raw view) -
    • Good CPU utilization: ~250k cores
      • ~60k cores for Analysis
      • ~190k cores for Production (note: CMS HLT back in operation ~25k cores)
    • No major issues otherwise

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • CNAF: Ticket Atlas about files with size 0 (GGUS:141996) Files being identified, in progress.
  • EGI: NC
  • FNAL: NC
  • IN2P3: NTR
  • JINR: NC
  • KISTI: NTR
  • KIT: NTR
  • NDGF: NC
  • NL-T1: Surfsara: configuration problem with the tape backend, last Friday 5th July. It was fixed around 17:20 CEST
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL:
    • Upgrading Production FTS tommorow (7/8/19) to 3.9.1 ( with DB schema update) and relevant gfal* and daviX* packages. NOT adding "scitoken" or "xrootd" packages (so still unavailable for TPC).
    • Batch Capacity nominal again after intervention for CMS.
  • TRIUMF: NTR

  • CERN computing services:
    • OTG:0051185 - Hypervisor reboot schedule to mitigate MDS Intel CPU vulnerabilities
      • Will proceed by availability zone, from 18th July. Please watch OTG for updates.
  • CERN storage services:
    • the CASTOR nameserver intervention originally planned for Thu Jul 18th has been shifted to Tue 23rd (all day). OTG0051004 and GocDB have been changed (original announce 2019-06-24, shifted 2019-07-04)
    • CASTOR: Short service unavailability this weekend due to network issue on Sun Jul 7th evening. As a consequence, some sessions were disconnected on node reboot.
    • CMS: EOSCMS was unreachable for ~10-15 minutes between 11:45 and 12:00 on Thu Jul 4th. We have gathered the needed information in order to understand the root cause and we will investigate this with the developers. OTG0051144
  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Draft availability report for June 2019 sent around
  • MW Officer: NC
  • Networks: NC
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2019-08-05 - BrianDavies
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback